Parallel to y-1=0 with y-intercept (0,-1) is the answer y=1 or y= -1, how do I arrive at the answer? thanks!
Answer by venugopalramana(3286)

Parallel to y-1=0 with y-intercept (0,-1) is the answer y=1 or y= -1, how do I arrive at the answer? thanks!
A PARALLEL LINE WILL HAVE SAME X&Y TERMS WITH A DIFFERENT CONSTANT
HENCE EQN OF PARALLEL TO
Y-1=0
WILL BE OF THE TYPE
Y-K=0
NOW Y INTERCEPT IS -1..THAT IS IF WE PUT X=0 , WE SHOULD GET Y=-1
WE HAVE Y-K=0
Y=K
HENCE K=-1
EQN. IS Y-(-1)=0
Y+1=0
OR...Y=-1